Find the center of mass for the region bounded by y=4-x^2 that is in the first quadrant.


1
Expert's answer
2021-04-14T12:36:32-0400

the region whose center of mass is desired is the region bounded by the curves "y=4-x^2", the positive x-axis and positive y-axis

since we are not given the mass density relation "\rho(x,y)" , we may take it uniform over the region.

thus we have

mass, "m= \int ^{2}_{0}\rho[4-x^2]dx= \rho [4x-\frac {x^2}{2}]^{2}_{0}"

"=6\rho" units of mass

moments about x=axis,"m_x= \frac 12 \rho \int ^{2}_{0}(4-x^2)^2dx=\frac12 \rho\int^{2}_{0}(16-8x^2+x^4)dx"

"=\frac 12\rho[16x-\frac83x^3+\frac 15x^5]^{2}_{0}=\frac12 \rho(\frac{256}{15})"

"=\frac{128}{15}\rho" units of the mass times units of length

and

moments about y-axis

"m_y= \rho\int^{2}_{0}x(4-x^2)dx= \rho\int^{2}_{0}(4x-x^3)dx"

"=\rho[2x^2-\frac {x^4}{ 4}]^{2}_{0}= 4\rho" units of mass times units of length

the center of mass "(\bar x, \bar y)" is such that "\bar x= \frac {m_y}{m} = \frac {4\rho}{6\rho}= \frac 23, \bar y = \frac {m_x}{m}= \frac {\frac {128\rho}{15}}{6\rho}=\frac {64}{45}"

"\implies (\frac 23, \frac {64}{45})" is the center of mass "y=4-x^2" in the first quadrant